Ms. Jackson let her students know that their final exam contained three groups of questions:
- 16 multiple choice worth 1.5 points each
- 16 short answer worth 2.5 points each
- 3 essay one worth 12 points each
A. What percent of the final exam is each group worth?
B. Beatrice has a semester average of 89 before she takes the final. She knows that the final is worth 25% of her total semester average. What is the greatest number of questions Beatrice can miss and still raise her final semester average to at least a 90? Explain how you arrived at your answer

Answer by richwmiller(17219)

A)
16*1.5+16*2.5+3*12=100 points for the exam.
multiple choice is 16*1.5 =24 percent of the exam
short answer is 16*2.5 =40 percent of the exam
essay=3*12=36 percent of the exam
B)
.25*x+.75*89=90
x=93
She needs to get a 93 on the exam.
So she can't lose more than 7 points to get her 90 average.